Assume, at MODUL university a representative sample of n=250 was drawn, 200 students were female.




What does this tell us about the percentage of female students studying at this university (the population)? (Hint: Calculate a 95%-confidence interval for your percentage prediction - namely the percentage value in the population.) (1p.)




Repeat the computation of the 95%-confidence interval with the following sample sizes: n=100, n=1500, and n=12000, each time assuming the same percentages as in 3). What happens and why? (1p.)





1
Expert's answer
2021-11-18T07:12:48-0500

"z=\frac{p-p_0}{\sqrt{\frac{p_0(1-p_0)}{n}}}"


for n= 250 percentage of female students:

"p_0=200\/250=4\/5=0.8"


critical values for 95%-confidence interval:

"z=\pm 1.96"


"-1.96<\frac{p-0.8}{0.05}<1.96"


"0.70<p_0<0.90"


for n= 100:

"-1.96<\frac{p-0.8}{0.04}<1.96"


"0.72<p_0<0.88"


for n= 1500:

"-1.96<\frac{p-0.8}{0.01}<1.96"


"0.78<p_0<0.82"


for n= 12000:

"-1.96<\frac{p-0.8}{0.004}<1.96"


"0.793<p_0<0.807"


Increasing the sample size decreases the width of confidence intervals, because it decreases the standard error.